# A Survey on Point-of-Interest Recommendation: Models, Architectures, and Security

### v1

@@ IV-B1 Long Short-Term Memory
LSTM networks have become indispensable for POI recommendation [80, 52, 81, 83, 84], effectively modeling the sequential dependencies inherent in user check-in data to predict future preferences. Earlier studies [80, 52, 81] adopt LSTM to model POI sequences. Zhang et al. [80] propose a novel approach to learning user preferences by introducing the Discrete Ranking-based Matrix Factorization (DRMF) algorithm. DRMF leverages pairwise user preferences, framing the learning process as a binary quadratic programming problem to derive binary codes. To address the challenges of non-convexity and binary constraints inherent in this formulation, they introduce a self-paced learning strategy. This strategy gradually incorporates pairwise preferences, starting with simpler relationships and progressing to more complex ones, thereby enhancing the optimization process. Nevertheless, Zhao et al. [52] address the limitations of traditional RNNs in capturing the crucial influence of time and distance between consecutive check-ins. Their proposed STLSTM model integrates spatio-temporal context, short-term spatio-temporal gating, and long-term spatio-temporal gating. Another earlier study [81] introduces an approach that incorporates temporal information to understand the semantic connections between user actions. This approach leverages a semantic correlational graph to calculate semantic sequential correlation, factoring in time intervals. It then integrates this semantic information into a modified LSTM framework, incorporating two additional semantic gates. This enhanced model captures users’ sequential behaviors and long- and short-term interests within a semantic context. Finally, this approach proposes user clustering to improve recommendation accuracy by grouping users based on their semantic preferences.
Recent studies [84, 87] based on LSTM mainly aim to capture dynamic dependencies in long-term POI sequences. Yang et al. [84] tackle the challenge of effectively capturing the dynamic interplay between a user’s long-term preferences and their more immediate, context-dependent interests, namely AMACF. It achieves this by dynamically weighting input sequences and seamlessly fusing long- and short-term signals. DLAN [87] incorporates a multi-head attention module that effectively combines first-order and higher-order neighborhood information within user check-in trajectories. This parallel approach addresses the limitations of RNN-based methods, which struggle to establish long-term dependencies between sequences. Furthermore, DLAN integrates a user similarity weighting layer to quantify the influence of social relationships between users on the target user’s preferences. This layer leverages social connections to improve the accuracy of recommendations by incorporating insights from similar users.
Beyond these examples, other studies such as Zhong et al. [82] and Zhang et al. [83] also explore innovative ways to leverage LSTMs for POI recommendation, further demonstrating the versatility and effectiveness of this approach. These studies highlight the ongoing evolution of LSTM-based methods for POI recommendation. Researchers continue to develop sophisticated techniques to capture the complexities of user behavior, ultimately leading to more accurate, personalized, and context-aware recommendations.
@@ IV-B3 Graph Neural Networks
Graph neural networks (GNNs) [159] provide a powerful framework for learning representations from graph-structured data. GNNs operate through a message-passing mechanism, where each node updates its representation by aggregating information from its neighboring nodes and then combining these aggregated messages with its current representation. An illustration of the GNN architecture is presented in Fig. 4 (a).
Consider a graph $G=(V,E)$, where $V$ is the set of nodes and $E$ is the set of edges. Each node $v\in V$ is associated with a feature vector $h^{\prime}_{v}$. GNNs utilize the structure of the graph, along with these node features $h^{\prime}_{v}$, to learn a representation vector for each node, $h_{v}$, or for the entire graph, $h_{G}$. Modern GNN architectures typically employ a neighborhood aggregation strategy, where a node’s representation is iteratively refined by aggregating the representations of its neighbors [160]. After $k$ iterations of aggregation, the node’s representation captures the structural information within its $k$-hop neighborhood. The $k$-th layer of a GNN is shown as:
@@
where $h_{v}^{(k)}$ is the feature vector of node $v$ at the $k$-th iteration, with $h_{v}^{(0)}$ initialized to $X_{v}$. The set $\mathcal{N}(v)$ denotes the nodes adjacent to $v$. The functions $\mathsf{Aggregate}^{(k)}(\cdot)$ and $\mathsf{Combine}^{(k)}(\cdot)$ are crucial to the design of GNNs.
GNNs have emerged as a powerful tool for POI recommendation [117, 20, 21, 118, 120, 123, 124], effectively capturing complex relationships within location-based data. Earlier studies [117, 20, 21, 118, 120] mainly emphasize capturing spatial and temporal dynamics. Specifically, Han et al. [117] focuses on capturing both spatial and temporal dynamics within user check-in data. First, it constructs a multigraph to represent various contextual factors and their relationships. Then, it employs a time-aware sampling strategy to capture the evolving influence of neighboring nodes. Finally, it learns node representations that change over time, reflecting the dynamic nature of user preferences and POI popularity. Another earlier study [20] proposes a novel next POI recommendation model called DynaPosGNN, which incorporates specific arrival times into offline user activities. Unlike traditional methods that solely rely on user check-in history, DynaPosGNN analyzes the correlation between arrival time and two spatial dynamic graphs: the “User-POI graph” and the “POI-POI graph”. By integrating temporal information, DynaPosGNN provides a more nuanced understanding of user behavior, considering both the locations visited and the timing of those visits, leading to potentially more accurate recommendations. In comparison, ASGNN [21] incorporates a personalized hierarchical attention network to exploit complex correlations between users and POIs in check-in sequences. This allows the model to capture both long-term and short-term user preferences, providing a more discriminative understanding of user behavior. Finally, ASGNN leverages these learned user preferences to predict the next POI, taking into account both long-term and short-term behavioral patterns. Zhang et al. [118] introduce GNN-POI, a POI recommendation framework that leverages the power of GNNs to learn node representations from both node information and topological structure, thus enhancing POI recommendation accuracy. GNNs have demonstrated exceptional capacity for modeling complex relationships within graph structures, making them well-suited for capturing the intricate connections between users and POIs in LBSNs.
@@
Recent studies [124, 123] based on GNN focus on adopting GNN to capture side information. Liu et al. [123] propose an approach to learning similar users’ POI transfer preferences using a session-aware GNN. This method aims to understand how users’ preferences for POIs evolve over time and how those preferences can be transferred to similar users. By contrast, HKGNN [124] leverages the power of hypergraphs and side information to enhance POI recommendations. The key details include utilizing a hypergraph to represent the complex relationships between users, POIs, and other relevant entities, employing a hypergraph neural network to learn from the rich structural information encoded in the HKG, and incorporating side information, such as POI categories and user demographics, to mitigate data sparsity and enhance recommendation accuracy.
@@ IV-C Self-Supervised Learning-Based
Self-supervised learning (SSL) [42] is a versatile framework that leverages unsupervised data by creating surrogate tasks, known as pretext tasks, to help models learn meaningful representations. These tasks are specifically designed to guide the learning process towards capturing useful features from the data. SSL methods are generally categorized into two approaches: generative and contrastive [161]. Generative SSL aims to create or reconstruct data, focusing on learning detailed representations, while contrastive SSL emphasizes comparing data samples by bringing similar pairs closer in the latent space and pushing dissimilar pairs apart.
The SSL process typically begins with defining a pretext task, which does not require labeled data but helps in learning valuable representations. In contrastive learning, the pretext task might involve differentiating between augmented versions of the same image (positive pairs) and different images (negative pairs). The model is trained by minimizing the following loss function [162]:
@@
where $\mathbf{1}_{[k\neq i]}\in\{0,1\}$ is an indicator function that equals 1 only if $k\neq i$, $\mathsf{pair}_{i}$ and $\mathsf{pair}_{j}$ are the feature representations of a positive pair, $\mathsf{Sim}(\cdot,\cdot)$ denotes the cosine similarity, $\tau$ is a temperature scaling parameter, and $N$ represents the number of data points.
@@
Self-supervised learning has emerged as a promising paradigm in POI recommendation [163]. By leveraging patterns and structures within data, self-supervised techniques enable models to learn meaningful representations without explicit labels. In POI recommendation, these methods capture latent user preferences and spatial contexts from unlabeled data, enhancing personalized recommendations. Self-supervised POI recommendation systems [125, 126, 128, 129, 44, 131, 132] can be divided into two categories. The first category [125, 126, 128] focuses on creating augmented views through human design. Zhou et al. [125] introduce a Self-supervised Mobility Learning (SML) framework to encode human mobility semantics and enhance location-based tasks. SML addresses challenges posed by sparse and noisy mobility data by leveraging spatio-temporal contexts and augmented traces. SML generates contrastive views via traditional ways like sampling trajectory points. Another work [126] introduces a paradigm that contrastively learns trajectory representations from geolocation and temporal transitions via random sampling POI points to obtain contrastive views, enhancing the model’s understanding of mobility patterns. At its core is a Graph-enhanced Self-attentive layer, which uncovers transitional dependencies among POIs and captures temporal interests. Lastly, SelfTrip [128] focuses on trip recommendation using a two-step contrastive learning mechanism to learn nuanced representations of POIs and itineraries based on generated views via random walk, further enhanced by four innovative trip augmentation methods.
@@
Nonetheless, recent studies [129, 44, 131, 132] on the self-supervised paradigm focus on creating augmented views automatically. While many existing studies prioritize final POI recommendation performance, they often overlook the crucial connection between POI sequences and contextual information. This can lead to limited data representations and less accurate recommendations. To address this, Jiang et al. [129] propose LSPSL, a unified attention framework for next POI recommendation that leverages self-supervised learning to model users’ long- and short-term preferences. LSPSL utilizes a self-attention network and two self-supervised optimization objectives to deeply explore the inherent relationships between POI sequences and contextual information and generate contrastive views automately during a pre-training phase. This process strengthens the data representations, allowing the model to better understand the nuances of user behavior. GSBPL [44] has been introduced for the domain of next-POI recommendation. GSBPL utilizes a pair of graph data augmentation techniques based on modeling user behaviour patterns via GNNs, to create augmented trajectory graphs, enabling the effective modeling of implicit behavioral patterns. To acquire high-order representations of trajectory graphs, GSBPL incorporates a graph preference representation encoder that utilizes geographical and social context. By applying contrastive learning, GSBPL captures implicit behavior patterns within these representations. To address the limitations of existing approaches that neglect the complex higher-order relationships between user-POI dynamics over time and fail to account for the varying importance of POIs, leading to an inability to effectively capture geographical influence imbalances, Fu et al. [132] introduce SLS-REC, a novel self-supervised long- and short-term model for POI recommendation. SLS-REC leverages a spatio-temporal attention hypergraph neural network to capture the spatial dependencies and temporal evolution of the short-term dynamic interests of users and creates contrastive views automately.
IV-D Generative Model-Based
@@ IV-D3 Variational Autoencoder
TABLE II: Model comparison in terms of effectiveness, efficiency, scalability and explainability
@@
Model Effectiveness Moderate High High High
Inference Efficiency High Varies Varies Low
Training Scalability High Moderate Moderate Low
Result Explainability Moderate Low Moderate High
IV-E Discussion of Models
In this part, we further compare the discussed model types: Latent Factor Models (LFMs), Classic Neural Networks (NNs), Self-Supervised Learning (SSL), and Generative Models (Gen-Models) from the aspects of model effectiveness, inference efficiency, training scalability, result explainability, shown in Table II. The comparison analysis is shown as:
Model effectiveness:
  • LFMs: LFMS moderately captures user-POI interactions using latent features for comparable recommendations.
  • NNs: NNs potentially achieve higher effectiveness by learning complex non-linear relationships between users and POIs, leading to more nuanced recommendations.
  • SSL: The effectiveness of SSL can be high, particularly when labeled data is scarce, by learning robust representations from unlabeled data.
  • Gen-Models: Gen-Models obtain high potential effectiveness due to their ability to generate diverse POI recommendations based on learned data distributions.
Inference efficiency:
  • LFMs: LFMS have high prediction efficiency, especially for smaller datasets, resulting in small prediction time.
  • NNs: Efficiency of NNs varies greatly depending on the network architecture and size; larger, more complex networks are less efficient.
  • SSL: Efficiency depends on the specific method employed; some approaches are more efficient than others.
  • Gen-Models: The inference efficiency of Gen-Models lower than other kinds of models due to their high model complexity and large model volume.
Training scalability:
  • LFMs: Scalability of training is higher than other models due to its simple model structure.
  • NNs: Training scalability is moderate; it’s limited only when the computational resources required for training complex networks become insufficient.
  • SSL: SSL often exhibits similar training scalability as NNs, especially methods using simpler architectures.
  • Gen-Models: Scalability is a major challenge; training scalability becomes computationally expensive with increasing data size.
Result explainability:
  • LFMs: Results of LFMs offer moderate explainability; latent factors provide some insights into user preferences.
  • NNs: Explainability is generally low; understanding results of complex neural networks is difficult.
  • SSL: Explainability is moderate; interpreting the learned representations requires further analysis.
  • Gen-Models: Explainability is higher than results of other models; understanding generated recommendation results often necessitates analyzing the latent space.
@@ V-A Centralized Architecture
Centralized architecture serves as a robust, scalable, and cost-effective foundation for building and deploying recommendation systems. In this model, centralized data centers and servers provide a consolidated platform for managing large-scale computational resources. By centralizing these resources, enterprises can swiftly respond to changes in demand, reduce overhead, and streamline system maintenance. A key benefit of this architecture is its ability to centralize data processing and storage, allowing for more efficient management of data flow and integration across systems. The flexibility of centralized architecture can be expressed through the following equation:
@@
This equation highlights how centralized systems optimize scalability, cost-effectiveness, and resource availability, enabling rapid adaptation without requiring extensive upfront hardware investments. We explore studies [51, 50, 88, 139, 53] on POI recommendation systems developed using centralized architecture. For instance, Wang et al. [51] leverage centralized computing to model geographical influences in POI recommendations with GeoIE. Centralized servers enable real-time processing of geographical data, enhancing the efficiency of recommendations by factoring in geo-influence, geo-susceptibility, and physical distance. Similarly, Yin et al. [50] developed SH-CDL, where centralized architecture plays a crucial role in managing the fusion of diverse POI features and user preferences across multiple locations. The centralization of data allows SH-CDL to more effectively model both individual and public preferences by leveraging collective data from a large number of users. Furthermore, Wang et al. [53] introduced AGRAN, which benefits from centralized data integration to create adaptive POI graphs and fuse POI representations with dynamic spatial-temporal data, ensuring real-time responsiveness to evolving user preferences. These studies underscore the advantages of centralized architecture in POI recommendation systems, emphasizing enhanced efficiency and scalability. However, as the complexity of these systems increases, continuous innovation in defense mechanisms is essential to safeguard data integrity and user privacy within centralized frameworks.
V-B Decentralized-based Architecture

### v2

@@ IV-B1 Long Short-Term Memory
LSTM networks have become indispensable for POI recommendation [56, 153, 57, 59, 60], effectively modeling the sequential dependencies inherent in user check-in data to predict future preferences. Earlier studies [56, 153, 57] adopt LSTMs to model POI sequences. Zhao et al. [153] address the limitations of traditional RNNs in capturing the crucial influence of time and distance between consecutive check-ins. Their proposed STLSTM model integrates spatio-temporal context, short-term spatio-temporal gating, and long-term spatio-temporal gating. Another study by Zhan et al. [57] incorporates temporal information to understand the semantic relationships between user actions. They leverage a semantic correlational graph to calculate semantic sequential correlation, factoring in time intervals, and integrate this information into a modified LSTM framework with two additional semantic gates. This enhanced model captures users’ sequential behaviors and long- and short-term interests within a semantic context. It also introduces user clustering to improve recommendation accuracy by grouping users based on their semantic preferences.
Recent studies [60, 63] based on LSTM mainly aim to capture dynamic dependencies in long-term POI sequences. Yang et al. [60] tackle the challenge of effectively capturing the dynamic interplay between a user’s long-term preferences and their more immediate, context-dependent interests with AMACF. It achieves this by dynamically weighting input sequences and seamlessly fusing long- and short-term signals. DLAN [63] incorporates a multi-head attention module that effectively combines first-order and higher-order neighborhood information within user check-in trajectories. This parallel approach addresses the limitations of RNN-based methods, which struggle to establish long-term dependencies between sequences. Furthermore, DLAN integrates a user similarity weighting layer to quantify the influence of social relationships between users on the target user’s preferences, leveraging social connections to improve the accuracy of recommendations.
Beyond these studies, others such as Zhong et al. [58] and Zhang et al. [59] explore innovative ways to leverage LSTMs for POI recommendation, further demonstrating the versatility and effectiveness of this approach. These studies highlight the ongoing evolution of LSTM-based methods for POI recommendation, as researchers continue to develop sophisticated techniques to capture the complexities of user behavior, ultimately leading to more accurate, personalized, and context-aware recommendations.
@@ IV-B3 Graph Neural Networks
Graph neural networks (GNNs) [160] provide a powerful framework for learning representations from graph-structured data. GNNs operate through a message-passing mechanism, where each node updates its representation by aggregating information from its neighboring nodes and then combining these aggregated messages with its current representation. An illustration of the GNN architecture is presented in Fig. 4 (a). Consider a graph $G=(V,E)$, where $V$ is the set of nodes and $E$ is the set of edges. Each node $v\in V$ is associated with a feature vector $h^{\prime}_{v}$. GNNs utilize the structure of the graph along with these node features $h^{\prime}_{v}$ to learn a representation vector for each node, $h_{v}$, or for the entire graph, $h_{G}$. Modern GNN architectures typically employ a neighborhood aggregation strategy, where a node’s representation is iteratively refined by aggregating the representations of its neighbors [161]. After $k$ iterations of aggregation, the node’s representation captures the structural information within its $k$-hop neighborhood. The $k$-th layer of a GNN is shown as:
@@
where $h_{v}^{(k)}$ is the feature vector of node $v$ at the $k$-th iteration, with $h_{v}^{(0)}$ initialized to $X_{v}$. The set $\mathcal{N}(v)$ denotes the neighbors of node $v$. The functions $\mathsf{Aggregate}^{(k)}(\cdot)$ and $\mathsf{Combine}^{(k)}(\cdot)$ are crucial components of GNN design.
GNNs have emerged as a powerful tool for POI recommendation [94, 21, 22, 95, 97, 100, 101], effectively capturing complex relationships within location-based data. Earlier studies [94, 21, 22, 95, 97] focus on capturing spatial and temporal dynamics. Specifically, Han et al. [94] captures both spatial and temporal dynamics in user check-in data. It constructs a multigraph to represent contextual factors and their relationships, employs a time-aware sampling strategy to account for the evolving influence of neighboring nodes, and learns time-dependent node representations to reflect changing user preferences and POI popularity. Another study [21] introduces DynaPosGNN, a next POI recommendation model that integrates arrival times into user activities. Unlike traditional methods that rely solely on check-in history, DynaPosGNN analyzes the correlation between arrival time and two spatial dynamic graphs—the ”User-POI graph” and the ”POI-POI graph”—providing a more nuanced understanding of user behavior. ASGNN [22] incorporates a personalized hierarchical attention network to capture complex correlations between users and POIs in check-in sequences. This enables the model to capture both long-term and short-term user preferences, enhancing its ability to predict the next POI. Finally, Zhang et al. [95] introduce GNN-POI, a POI recommendation framework that utilizes GNNs to learn node representations from both node information and topological structure, improving POI recommendation accuracy. GNNs’ exceptional capacity to model complex relationships makes them well-suited for capturing the intricate connections between users and POIs in LBSNs.
@@
Recent studies [101, 100] based on GNN focus on using GNNs to capture side information. Liu et al. [100] propose a session-aware GNN to learn POI transfer preferences from similar users. This approach aims to understand how users’ POI preferences evolve over time and how these preferences can be transferred to similar users. In contrast, HKGNN [101] utilizes hypergraphs and side information to enhance POI recommendations. It employs a hypergraph to represent complex relationships between users, POIs, and other entities, a hypergraph neural network to learn from the structural information encoded in the hypergraph, and side information such as POI categories and user demographics to address data sparsity and improve recommendation accuracy.
@@ IV-C Self-Supervised Learning-Based
Self-supervised learning (SSL) [149] is a versatile framework that leverages unsupervised data by creating surrogate tasks, known as pretext tasks, to help models learn meaningful representations. These tasks are designed to guide the learning process towards capturing useful features from the data. SSL methods are generally categorized into two approaches: generative and contrastive [162]. Generative SSL aims to create or reconstruct data, focusing on learning detailed representations, while contrastive SSL emphasizes comparing data samples by bringing similar pairs closer in the latent space and pushing dissimilar pairs apart.
The SSL process typically begins by defining a pretext task, which does not require labeled data but helps in learning valuable representations. In contrastive learning, the pretext task involves differentiating between augmented versions of the same image (positive pairs) and different images (negative pairs). The model is trained by minimizing the following loss function [163]:
@@
where $\mathbf{1}_{[k\neq i]}\in\{0,1\}$ is an indicator function that equals 1 only if $k\neq i$, $\mathsf{pair}_{i}$ and $\mathsf{pair}_{j}$ are the feature representations of a positive pair, $\mathsf{Sim}(\cdot,\cdot)$ denotes cosine similarity, $\tau$ is a temperature scaling parameter, and $N$ represents the number of data points.
@@
SSL has emerged as a promising paradigm in POI recommendation [164]. By leveraging patterns and structures within data, self-supervised techniques enable models to learn meaningful representations without explicit labels. In POI recommendation, these methods capture latent user preferences and spatial contexts from unlabeled data, enhancing personalized recommendations. Self-supervised POI recommendation systems [102, 103, 105, 107, 106, 109, 110] can be divided into two categories. The first focuses on creating augmented views through human design [102, 103, 105]. Zhou et al. [102] introduced a Self-supervised Mobility Learning (SML) framework to enhance location-based tasks by encoding human mobility semantics. SML addresses challenges posed by sparse and noisy mobility data by leveraging spatio-temporal contexts and augmented traces. It generates contrastive views through methods such as sampling trajectory points. Another work [103] introduces a paradigm that learns trajectory representations via random sampling of POI points, enhancing the model’s understanding of mobility patterns. A Graph-enhanced Self-attentive layer uncovers transitional dependencies among POIs and captures temporal interests. Lastly, SelfTrip [105] focuses on trip recommendation using a two-step contrastive learning mechanism to learn nuanced representations of POIs and itineraries based on augmented views from random walks, further enhanced by four innovative trip augmentation methods.
@@
Recent studies [107, 106, 109, 110] focus on automatically creating augmented views. While many studies prioritize POI recommendation performance, they often overlook the connection between POI sequences and contextual information, leading to limited data representations and less accurate recommendations. To address this, Jiang et al. [107] propose LSPSL, a unified attention framework for next-POI recommendation that leverages self-supervised learning to model long- and short-term preferences. LSPSL uses a self-attention network and two self-supervised optimization objectives to explore relationships between POI sequences and contextual information, generating contrastive views automatically during pre-training. GSBPL [106] introduces graph-based data augmentation techniques for next-POI recommendation, modeling user behavior patterns via GNNs to create augmented trajectory graphs. By applying contrastive learning, GSBPL captures implicit behavioral patterns. To address limitations in capturing higher-order relationships and the varying importance of POIs, Fu et al. [110] introduce SLS-REC, a self-supervised model for POI recommendation. SLS-REC leverages a spatio-temporal attention hypergraph neural network to capture spatial dependencies and the temporal evolution of short-term dynamic user interests while automatically creating contrastive views.
IV-D Generative
@@ IV-D3 Variational Autoencoder
IV-E Discussion of Models
In this section, we compare the model types discussed earlier: Latent Factor Models (LFMs), Classic Neural Networks (NNs), Self-Supervised Learning (SSL), and Generative Models (Gen-Models), focusing on model effectiveness, inference efficiency, training scalability, and result explainability, as summarized in Table III.
TABLE III: Comparison of models based on effectiveness, efficiency, scalability, and explainability
@@ IV-E Discussion of Models
Effectiveness Moderate High High High
Efficiency High Varies Varies Low
Scalability High Moderate Moderate Low
Explainability Moderate Low Moderate High
Firstly, in terms of model effectiveness, LFMs provide moderate performance by using latent features to generate reasonable recommendations, though with limited depth. NNs, on the other hand, excel at capturing complex relationships, leading to more nuanced recommendations. SSL is particularly effective when labeled data is scarce, as it learns robust representations from unlabeled data. Gen-Models offer high effectiveness, especially when diverse recommendations are needed, due to their generative nature.
Secondly, regarding inference efficiency, LFMs are highly efficient, particularly with smaller datasets. NNs’ efficiency depends on the network architecture, with larger models resulting in slower inference. The efficiency of SSL varies depending on the method used, with some being computationally expensive. Gen-Models tend to be less efficient due to their complexity and large parameter volumes, which result in longer inference times.
Additionally, in terms of training scalability, LFMs are highly scalable due to their simple structure, enabling efficient handling of large datasets. NNs and SSL have moderate scalability, as their training becomes computationally intensive with larger datasets. Gen-Models face significant scalability challenges, as training them becomes prohibitively expensive with the increasing data size.
Finally, in terms of result explainability, LFMs offer moderate transparency, providing some insight into user preferences through latent factors. NNs generally lack explainability due to their complex architectures, making it difficult to interpret their decisions. SSL provides moderate explainability, though interpreting the learned representations often requires further analysis. Gen-Models provide higher explainability than the other models, but understanding them fully necessitates exploring the latent space, which can be complex.
@@ V-A Centralized Architecture
Centralized architecture provides a scalable, cost-effective foundation for recommendation systems by using centralized data centers to manage computational resources. This model allows enterprises to respond quickly to demand changes, reduce overhead, and streamline maintenance. Centralizing data processing and storage enhances efficiency in managing data flow and system integration. The flexibility of centralized architecture is expressed by the following equation:
@@
This equation illustrates how centralized systems optimize scalability, cost-effectiveness, and resource availability with minimal upfront hardware investments. We explore studies [120, 121, 65, 122, 123] on POI recommendation systems developed using centralized architecture. For instance, Wang et al. [120] leveraged centralized computing to model geographical influences in POI recommendations with GeoIE. Centralized servers enable real-time processing of geographical data, enhancing the efficiency of recommendations by factoring in geo-influence, geo-susceptibility, and physical distance. Similarly, Yin et al. [121] developed SH-CDL, where centralized architecture plays a crucial role in managing the fusion of diverse POI features and user preferences across multiple locations. The centralization of data allows SH-CDL to more effectively model both individual and public preferences by leveraging collective data from a large number of users. Furthermore, Wang et al. [123] introduced AGRAN, which benefits from centralized data integration to create adaptive POI graphs and fuse POI representations with dynamic spatial-temporal data, ensuring real-time responsiveness to evolving user preferences. These studies underscore the advantages of centralized architecture in POI recommendation systems, emphasizing enhanced efficiency and scalability. However, as the complexity of these systems increases, continuous innovation in defense mechanisms is essential to safeguard data integrity and user privacy within centralized frameworks.
V-B Decentralized Architecture